OpenStudy (anonymous):

lim as x-> 5pi/2 of (sin^2x + 6sinx + 5)/sin^2x-1 Please explain HOW you do it...

OpenStudy (jamesj):

First I would simplify this expression and see what you can see.

OpenStudy (jamesj):

hint: factor in terms of sin x

OpenStudy (anonymous):

I got it to (sinx + 5)(sinx + 1) / (sinx+1)(sinx-1)

OpenStudy (anonymous):

but cancelling doesn't help

OpenStudy (jamesj):

right, now cancel one of these terms. And then try and take the limit.

OpenStudy (anonymous):

because sin(5pi/2) is 1 and 1-1 is 0..which makes the denominator 0

OpenStudy (jamesj):

yes, the denominator approaches zero from below. And the numerator is always positive. Hence the limit is improper and what is it?

OpenStudy (anonymous):

I don't understand

OpenStudy (jamesj):

The numerator is positive, the denominator is negative and approaches zero as x --> 5pi/2. Hence the limit must be ... what?

OpenStudy (anonymous):

negative infinity? but how do you know the denominator is negative?

OpenStudy (jamesj):

Yes, -infty. Because for all values of x, \( \sin x \leq 1 \) and in a neighborhood around x = 5pi/2--but not including 5pi/2, sin x is strictly less than 1. Therefore sin x - 1 < 0

OpenStudy (anonymous):

what is "\( \sin x \leq 1 \)"

OpenStudy (jamesj):

sin x is less than or equal to 1.

OpenStudy (anonymous):

ok and how do you know it's less than 0?

OpenStudy (jamesj):

because sin(5pi/2) = 1 but for any x near 5pi/2, sin(x) < 1. Hence for such values of x: sin(x) - 1 < 0

OpenStudy (jamesj):

Here's an alternative way. sin^2 x - 1 = -cos^2 x. Now take the limit of the whole original expression and you'll find the same answer.

OpenStudy (anonymous):

OH wait I think I understand

OpenStudy (anonymous):

at 5pi/2, on the unit circle, sin is at its highest

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

if you approach it from the left or right, it has to be less than 1 because it is a circle

OpenStudy (anonymous):

ok I understand how it is negative, why is it infinity?

OpenStudy (jamesj):

negative infinity. Because the numerator is positive, and the denominator is negative going to zero.

OpenStudy (anonymous):

yes I mean where does the infinity part come from

OpenStudy (anonymous):

i understand that the answer has to be negative

OpenStudy (jamesj):

What's the limit of \[ \lim_{x \to 0} \frac{17}{-x^2} \]

OpenStudy (anonymous):

lim x-> 0 of 17/-x^2 ? sorry, but for some reason your text is all messed up

OpenStudy (anonymous):

i think it's because the equation part of openstudy isn't working

OpenStudy (jamesj):

Yes, that limit is -infty

OpenStudy (anonymous):

yes because the numerator is constant and the bottom is smaller and smaller but still negative

OpenStudy (jamesj):

something's wrong with your browser then. But in any case, you see why that's the answer? The same thing is true here with your limit. lim x -> 5pi/2 of (sin x + 5)/(sin x - 1) The numerator is ALWAYS positive, but the denominator is negative and going to zero.

OpenStudy (anonymous):

is that problem supposed to be really tricky?

OpenStudy (anonymous):

haha I understand it but i feel like it's very obscure

OpenStudy (jamesj):

It tests your understanding. Try it the other way I proposed using sin^2 x - 1 = -cos^2 x You might find that a bit more intuitive. But in any case, one of the reason we do these problems is to work out the kinks in our conceptual understanding.

OpenStudy (anonymous):

yeah it really blew my mind when I finally understood it, thanks!